OpenStudy (anonymous):

write the expression as a simple exponential expression of the form a^b. -3^2/3^4

OpenStudy (jdoe0001):

\(\bf \cfrac{-3^2}{3^4} \implies \cfrac{-3 \times 3 }{3 \times 3\times 3\times 3}\)

OpenStudy (jdoe0001):

cancel like-terms

OpenStudy (anonymous):

-9/81?